Two cars approach an ice-covered intersection. One car, of mass 1.11*103 kg, is initially traveling north at 12.1 m/s. The other car, of mass 1.70*103 kg, is initially traveling east at 12.1 m/s. The cars reach the intersection at the same instant, collide, and move off coupled together. Find the velocity of the center of mass of the two-car system just after the collision.

Magnitude=

Directions = North of East
Answer #1

Here ,

let north be the positive y direction

east be the positive x direction

v1 = 12.1 j m/s

v2 = 12.1 i m/s

let the final speed is v

Using conseravtion of momentum

m1 * v1 + m2 * v2 = (m1 + m2) * v

12.1 * 1.11 *10^3 j + 1.7 *10^3 * 12.1 i = (1.11 *10^3 + 1.7 *10^3)* v

v = 4.78 j + 7.32 i

v= sqrt(7.32^2 + 4.78^2) at arctan(4.78/7.32)

v = 8.47 m/s at 33.14 degree

the magnitude of the velocity is 8.47 m/s is 33.14 degree

the magnitude of velocity os 8.47 m/s

direction of velocity is 33.14 degree north of east